what is the denisty in g/L STP if the gas in a 1.00L bulb weighs 0.672g of 25 celsius and 733.4 mmhg pressure? please show the work.
Answer #1

if we start with this equation
.. PV = nRT

and this one.. from definition of moles
.. n = mass / mw... . . (mw is my shorthand for molecular weight.. molar mass etc)

substitute
.. PV = (mass / mw) RT

rearranging
.. (mass / V) = mw x P / (RT)

since density = mass / Volume... and using ρ to represent density
.. ρ = mw x P / (RT)

rearranging
.. (mw x P) / (ρ x T) = R

now we all know that R is a constant so that ALL (mwP) / (ρT)'s must be equal
so that for 2 gases.. or 1 gas in 2 conditions
.. (mw1 x P1) / (ρ1 x T1) = R = (mw2 x P2) / (ρ2 x T2)

and since we're not changing gases.. mw1 = mw2 and that term drops out leaving
.. .. P1 / (ρ1 x T1) = P2 / (ρ2 x T2)

rearranging one more time
.. ρ2 = ρ1 x (P2 / P1) x (T1 / T2)

********
and the rest is straightforward..
.. ρ2 = (0.672g / L) x (760mmHg / 733.4mmHg) x (298.15K / 273.15K) = 0.760 g/L